True or False : 3Z = Z + Z + Z when Z is a matrix
1
Expert's answer
2021-10-05T15:50:12-0400

Answer: true


Using commutative property for matrix A+B=B+A,

and associative property for matrix A+(B+C)=(A+B)+C,

we get:

Z+Z=2Z

Z+Z+Z=Z+2Z=3Z